OpenStudy (moongazer):

could you show me how to find the quotient of this. (x^6+y^6 )÷(x+y)

OpenStudy (anonymous):

no need to divide \(x^6+y^6=x^6-y^6+2y^6 \)

OpenStudy (moongazer):

sorry, but what is that?

OpenStudy (moongazer):

it is asking for the quotient

OpenStudy (anonymous):

\( x^6+y^6=x^6-y^6+2y^6=(x^3-y^3)(x^3+y^3)+2y^6\\=(x-y)(x^2+xy+y^2)(x+y)(x^2-xy+y^2)+2y^6\\=(x+y)(x-y)(x^2+xy+y^2)(x^2-xy+y^2)+2y^6 \) ------------------------------ -- quotient re

OpenStudy (moongazer):

I understood the other parts. but why is it\[x ^{6} + y^{6}=x ^{6} - y^{6}+2y ^{6}\]

OpenStudy (anonymous):

because 2y^6-y^6=y^6

OpenStudy (moongazer):

ohh ok

OpenStudy (moongazer):

What should be my final answer?

OpenStudy (anonymous):

finall answer is up there in my second reply